# Weak-Lensing Shear-Selected Galaxy Clusters from the Hyper Suprime-Cam Subaru Strategic Program: I. Cluster Catalog, Selection Function and Mass–Observable Relation

### v1

@@ 4.3.2 Consistency in the Modeling Framework
To ensure the self-consistency and the correctness of our modeling framework, we further compare the predicted number counts from eq. (LABEL:eq:number_count_model) with the number counts obtained from a mock observation and from the real observed catalog.
First, we test whether our modeling framework can recover the number counts of a mock sample generated using a set of known cosmological and nuisance parameters. The mock catalog is created as a Poisson realization of the halo mass function (Bocquet et al., 2016) up to redshift 2 over an area of $\Omega_{\mathrm{survey}}=46.37,\deg^{2}$. Each mock cluster is assigned a weak-lensing mass, $M_{\mathrm{WL}}$, with respect to its true mass following a log-normal distribution, $\ln M_{\mathrm{WL}}\sim\mathcal{N}(\ln M,0.18^{2})$. Similarly, the concentration of each mock cluster is randomly drawn from a log-normal distribution with a mean value predicted by Diemer & Joyce (2019) and a scatter of $\sigma_{\ln c}=0.30$. Following steps (ii)-(iv) in Sec.,3.2, these mock clusters are then injected into the XMM subfield of the HSC-Y1 shape catalog, with the detected clusters forming a mock sample of shear-selected peaks.
The number counts of the mock sample is shown as orange circles in Fig. 7. The excenllent agreement between the number counts from a mock observations and those predicted by eq. (LABEL:eq:number_count_model) indicates that the modeling framework we develop is self-consistent. Moreover, we compare these results directly with the number counts of shear-selected clusters observed in this field (green squares). The good agreement (at a level of $\lesssim 1\sigma$) seen in Fig. 7 suggests that the modeling framework provides a good description of the data.
@@
Figure 8: Validation test on the shape of our predicted number counts of weak lensing peaks’ optical counterparts as a function of their optical richness given the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$ (see step (v) in Sec. 3.2). These predictions are compared with the number counts of the closest CAMIRA counterpart for weak lensing peaks found in the HSC-Y1 XMM field with a signal-to-noise ratio larger than 4.7 (green squares). A statistically more constraining result using the entire HSC-Y3 field obtained after unblinding the cosmological analysis in Chiu et al. (2024) is shown in Fig. 9 in Appendix A.
@@ 4.3.3 Comparison with Cross-matched Optical Clusters
Similar to eq. (LABEL:eq:number_count_model), we can write the number counts of optical counterparts as a function of their richness as
@@
We make use of the richness–mass scaling relation from Murata et al. (2019) to evaluate $P\left(N_{\mathrm{mem}}|M,z\right)$. Together with the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived from the injection simulations, we can compute the predicted richness distribution of the shear-selected clusters. When evaluating eq. (32), we note that we restrict the redshift integral to be $z\geq 0.1$, as we compare the richness distribution with that of the optical counterparts from CAMIRA which only covers $z\geq 0.1$.
@@
Fig. 8 shows the comparison between the observed richness distribution and the theoretical predictions using the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$ (see step (v) in Sec. 3.2). This plot is generated with the cosmological parameters fixed to those measured by Planck. The most distinct feature of changing the parameter $\Delta_{s}$ is revealed by the overall shape of $\mathrm{d}N/\mathrm{d}N_{\mathrm{mem}}$: The smaller (larger) value of $\Delta_{s}$ results in a decreasing (increasing) number of clusters at the low-richness end. This is expected, given that the more low-mass systems would be up-scattered into the shear-selected sample with an increasing $\Delta_{s}$. We see that the observed richness distribution in the HSC-Y1 XMM field (green squares) demonstrates agreement with the richness distributions predicted with $\Delta_{s}\sim 4.5$. A statistically more constraining result using the entire HSC-Y3 field obtained after unblinding the cosmological analysis in Chiu et al. (2024) is shown in Fig. 9 in Appendix A.
@@
We stress that the normalization and shape of $\mathrm{d}N/\mathrm{d}N_{\mathrm{mem}}$ are also affected by the cosmological parameters (e.g., $\Omega_{\mathrm{m}}$ and $\sigma_{8}$) and the scatter in both the richness–mass scaling relation and the weak-lensing mass bias. Therefore, we must simultaneously constrain the parameter $\Delta_{s}$ together with all the cosmological and nuisance parameters in a self-calibrating way. Moreover, because the optical counterparts shown in Fig. 8 are determined by only the matching in the sky coordinate, we expect some falsely matched counterparts at the low-richness end compared to that obtained from the more advanced matching including both the cluster redshift and richness (See Appendix A). It is feasible but challenging to include the richness distribution of the correctly matched counterparts into the data vector and self-consistently constrain the parameter $\Delta_{s}$. We therefore do not attempt to determine $\Delta_{s}$ here and leave it to future work. In Chiu et al. (2024), $\Delta_{s}$ will be constrained together with other parameters based on the number counts $N\left(\nu\right)$ only.
@@ 5 Conclusions
We have validated the weak-lensing scaling relation derived from our injection scheme is indeed insensitive to changes of cosmological parameters. We have derived three sets of scaling relations under three reference cosmologies (Table 1). The resulting weak-lensing peak number counts derived from these scaling relations have been shown to agree with each other excellently (Fig. 7). We have also compared these model predictions with the number counts of a mock sample obtained by directly injecting a realization of halo catalog in the weak-lensing mass maps. These tests have demonstrated the self-consistency of our modeling framework. The correctness of our model has been further verified by comparing our predictions with observed number counts of shear-selected clusters as a function of both weak-lensing signal-to-noise ratio and richness of the optical counterparts. These validation tests have been carried out using a small subfield of the HSC-Y1 shear catalog so that we still respect the blinded analysis.
@@ Acknowledgements
K.-F.C. acknowledges support from the Taiwan Think Global Education Trust Scholarship and the Taiwan Ministry of Education’s Government Scholarship to Study Abroad. K.-F.C. thanks the computing resources provided by the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan which are maintained by Dr. Bau-Ching Hsieh and others. K.-F.C. thanks the hospitality of Adrian Liu at McGill University where this work is carried out in part. I-Non Chiu acknowledges the support from the National Science and Technology Council in Taiwan (Grant NSTC 111-2112-M-006-037-MY3) and the computing resources provided by the National Center for High-Performance Computing (NCHC) in Taiwan. YTL acknowledges supports by the grant NSTC 112-2112-M-001-061. This work was supported by JSPS KAKENHI Grant Numbers JP20H05856, JP22H01260, JP22K21349.
@@
This paper makes use of software developed for Vera C. Rubin Observatory. We thank the Rubin Observatory for making their code available as free software at http://pipelines.lsst.io/. Numerical calculations are performed through the Python packages numpy (Oliphant, 2006–) and scipy (Virtanen et al., 2020). Many of the cosmological and astrophysical calculations in this work rely on the routines wrapped up in colossus (Diemer, 2018) and astropy (Astropy Collaboration et al., 2022). Plots are made available thanks to matplotlib (Hunter, 2007) and seaborn (Waskom, 2021).
@@ Appendix A Confidence level of Cross-matched Optical Counterparts
Murata et al. (2019) obtained constraints on the parameter $(A,B,B_{z},C_{z},\sigma_{0},q,q_{z},p_{z})$ by fitting to the HSC-Y1 shear data assuming a pivot mass $M_{\mathrm{pivot}}=3\times 10^{14}\,h^{-1}M_{\odot}$ and redshift $z_{\mathrm{pivot}}=0.6$. Here, we adopt the constraint derived under the Planck Collaboration et al. (2020) cosmology $(A,B,B_{z},C_{z},\sigma_{0},q,q_{z},p_{z})=(3.36,0.83,-0.20,3.51,0.19,-0.02,0.23,1.26)$. We note that this constraint is obtained for the optical clusters in the CAMIRA catalog with $N_{\mathrm{mem}}\geq 15$ and $0.1\leq z_{\mathrm{cl}}\leq 1.0$. Therefore, we only cross-match to CAMIRA clusters within these ranges. For the other terms in eq. (34), $P\left(\nu|M,z\right)$ is obtained following the treatment in Sec. 4.3
@@
Figure 9: Same as Fig. 8, number counts of weak lensing peaks’ optical counterparts as a function of their optical richness given the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$. These results are derived with $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ for HSC-Y3 after we unblind the cosmological analysis in Chiu et al. (2024). The green squares show the number counts of the closest CAMIRA counterpart for weak-lensing peaks in the entire HSC-Y3 footprint with a signal-to-noise ratio larger than $4.7$.

### v2

@@ 4.3.2 Consistency in the Modeling Framework
To ensure the self-consistency and the correctness of our modeling framework, we further compare the predicted number counts from Equation (LABEL:eq:number_count_model) with the number counts obtained from a mock observation and from the real observed catalog.
First, we test whether our modeling framework can recover the number counts of a mock sample generated using a set of known cosmological and nuisance parameters. The mock catalog is created as a Poisson realization of the halo mass function (Bocquet et al., 2016) up to redshift 2 over an area of $\Omega_{\mathrm{survey}}=46.37,\deg^{2}$. Each mock cluster is assigned a weak-lensing mass, $M_{\mathrm{WL}}$, with respect to its true mass following a log-normal distribution, $\ln M_{\mathrm{WL}}\sim\mathcal{N}(\ln M,0.18^{2})$. Similarly, the concentration of each mock cluster is randomly drawn from a log-normal distribution with a mean value predicted by Diemer & Joyce (2019) and a scatter of $\sigma_{\ln c}=0.30$. Following step (ii)-(iv) in Sec. 3.2, these mock clusters are then injected into the XMM subfield of the HSC-Y1 shape catalog, with the detected clusters forming a mock sample of shear-selected peaks.
The number counts of the mock sample is shown as orange circles in Figure 7. The excenllent agreement between the number counts from a mock observations and those predicted by Equation (LABEL:eq:number_count_model) indicates that the modeling framework we develop is self-consistent. Moreover, we compare these results directly with the number counts of shear-selected clusters observed in this field (green squares). The good agreement (at a level of $\lesssim 1\sigma$) seen in Figure 7 suggests that the modeling framework provides a good description of the data.
@@
Figure 8: Validation test on the shape of the predicted number counts of weak lensing peaks’ optical counterparts as a function of their optical richness given the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$ (see step (v) in Sec. 3.2). These predictions are compared with the number counts of the closest CAMIRA counterpart for weak lensing peaks found in the HSC-Y1 XMM field with a signal-to-noise ratio larger than 4.7 (green squares). A statistically more constraining result using the entire HSC-Y3 field obtained after unblinding the cosmological analysis in Chiu et al. (2024) is shown in Figure 9 in Appendix A.
@@ 4.3.3 Comparison with Cross-matched Optical Clusters
Similar to Equation (LABEL:eq:number_count_model), we can write the number counts of optical counterparts as a function of their richness as
@@
We make use of the richness–mass scaling relation from Murata et al. (2019) to evaluate $P\left(N_{\mathrm{mem}}|M,z\right)$. Together with the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived from the injection simulations, we can compute the predicted richness distribution of the shear-selected clusters. When evaluating Equation (32), we note that we restrict the redshift integral to be $z\geq 0.1$, as we compare the richness distribution with that of the optical counterparts from CAMIRA which only covers $z\geq 0.1$.
@@
Figure 8 shows the comparison between the observed richness distribution and the theoretical predictions using the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$ (see step (v) in Sec. 3.2). This plot is generated with the cosmological parameters fixed to those measured by Planck. The most distinct feature of changing the parameter $\Delta_{s}$ is revealed by the overall shape of $\mathrm{d}N/\mathrm{d}N_{\mathrm{mem}}$: The smaller (larger) value of $\Delta_{s}$ results in a decreasing (increasing) number of clusters at the low-richness end. This is expected, given that the more low-mass systems would be up-scattered into the shear-selected sample with an increasing $\Delta_{s}$. We see that the observed richness distribution in the HSC-Y1 XMM field (green squares) favors a higher value in $\Delta_{s}$. A statistically more constraining result using the entire HSC-Y3 field obtained after unblinding the cosmological analysis in Chiu et al. (2024) is shown in Figure 9 in Appendix A.
@@
We stress that the normalization and shape of $\mathrm{d}N/\mathrm{d}N_{\mathrm{mem}}$ are also affected by the cosmological parameters (e.g., $\Omega_{\mathrm{m}}$ and $\sigma_{8}$) and the scatter in both the richness–mass scaling relation and the weak-lensing mass bias. Therefore, we must simultaneously constrain the parameter $\Delta_{s}$ together with all the cosmological and nuisance parameters in a self-calibrating way. Moreover, because the optical counterparts shown in Figure 8 are determined by only the matching in the sky coordinate, we expect some falsely matched counterparts at the low-richness end compared to that obtained from the more advanced matching including both the cluster redshift and richness (See Appendix A). It is feasible but challenging to include the richness distribution of the correctly matched counterparts into the data vector and self-consistently constrain the parameter $\Delta_{s}$. We therefore do not attempt to determine $\Delta_{s}$ here and leave it to future work. In Chiu et al. (2024), $\Delta_{s}$ will be constrained together with other parameters based on the number counts $N\left(\nu\right)$ only.
4.4 Discussions
In previous studies, weak-lensing peaks are usually modeled either with an analytical halo model or with numerical simulations. An analytical framework models weak-lensing peaks as halos on top of a Gaussian random field, and the selection function is assumed to be a complementary error function (e.g. Kruse & Schneider, 1999, 2000; Bartelmann et al., 2002; Hamana et al., 2004; Fan et al., 2010; Miyazaki et al., 2018). These methods are similar to the “shape noise only” case in our analysis. While in some studies (Fan et al., 2010; Miyazaki et al., 2018), large-scale structures were also taken into account in addition to shape noise, it was still assumed that fluctuations due to large-scale structures are Gaussian. Therefore, these results are similar to our “shape noise only” case with slightly higher noise level. However, our injection simulation demonstrates that the fluctuation in the signal-to-noise ratio of a peak is non-Gaussian (see Figure 5 and 6). We attribute this to the non-linear growth of large-scale structures and the variation in shape noise due to the changing survey depth. As a result, we see in Figure 7 that the predicted number counts from the shape-noise-only case deviate from the injection simulation at lower signal-to-noise ratios as the impact of non-Gaussian fluctuations becomes stronger.
On the other hand, these non-Gaussian fluctuations can also be fully captured through numerical simulations (e.g. Hamana et al., 2004; Hennawi & Spergel, 2005; Wang et al., 2009; Kratochvil et al., 2010; Yang et al., 2011). However, N-body and ray-tracing simulations are expensive to conduct and are usually only carried out for a limited number of parameter sets. To partly circumvent this challenge, Lin & Kilbinger (2015) developed a “fast simulation” by randomly injecting analytical halo profiles into N-body simulations. Even though these randomly injected halos do not correlate with the underlying N-body simulations, Lin & Kilbinger (2015) showed that this semi-analytical approach can faithfully reproduce the peak abundance. The formalism developed in this work is similar in spirit but different in two major ways: 1) In our work, instead of injecting analytical halo profiles into N-body simulations, we inject these halos into the observed shear catalog. This allows us to estimate the line-of-sight scattering according to the observed underlying structures of our universe instead of relying on numerical prescriptions. Moreover, we calculate injected weak-lensing signal on individual galaxies which enables us to directly account for fluctuations due to shape noise. 2) Instead of drawing halos from a halo mass function to directly infer the number counts of weak lensing peaks, we uniformly sample halos across a large parameter space to focus exclusively on deriving the scaling relation and the selection function. The number counts are obtained by convolving our results with a halo mass function (see Equation (LABEL:eq:number_count_model)). By combining with the parametrization scheme introduced in Sec. 3.3, our injection simulation only needs to be performed once to capture the measurement uncertainties. Other systematic effects such as photometric redshift uncertainties and deviations from the spherical NFW profile can be easily incorporated during the modeling stage (see Sec. 4.3 and Chiu et al. 2024).
Even with full N-body and ray-tracing simulations, additional systematic effects related to cluster member galaxies and baryonic physics still need to be considered. The main systematic effects on peak statistics are dilution from cluster member galaxies, intrinsic alignment of satellite galaxies in cluster halos, and baryonic effects on the density profile. These systematic effects bias the peak signal-to-noise ratio and therefore affect cosmological constraints. Previous studies often made explicit corrections to these systematic effects that are around $10\%$ (Applegate et al., 2014; Melchior et al., 2015; Kacprzak et al., 2016). In our work, we adopt a conservative strategy instead to generate a contamination-free weak-lensing observable. Oguri et al. (2021) carefully examined different choices of filter shape and source galaxy redshift cut. It was shown that the combination of the filter shape adopted in this work and a source redshift cut of $0.7$ can effectively remove the contributions from cluster member galaxies. With this combination, the source galaxy number density around galaxy clusters in the region where the filter is non-zero matches that of the random field (see Fig. 9 in Oguri et al., 2021). This ensures that the weak-lensing maps we generate do not contain dilution from cluster member galaxies and intrinsic alignments of galaxies around the peak. Alignments of source galaxies with other structures along the line-of-sight are taken into account through the injection simulation as we inject signals into the observed shear catalog which already contains the effect of intrinsic alignments. Our adopted filter also removes contributions from cluster centers so that we are less sensitive to uncertainties in the modeling of cluster cores. The remaining uncertainties in halo density profiles are captured by a weak-lensing mass bias relation (Grandis et al., 2021; Chiu et al., 2022) that is self-consistently calibrated alongside our cosmological constraints (Chiu et al., 2024).
@@ 5 Conclusions
We have validated the weak-lensing scaling relation derived from our injection scheme is indeed insensitive to changes of cosmological parameters. We have derived three sets of scaling relations under three reference cosmologies (Table 1). The resulting weak-lensing peak number counts derived from these scaling relations have been shown to agree with each other excellently (Figure 7). We have also compared these model predictions with the number counts of a mock sample obtained by directly injecting a realization of halo catalog in the weak-lensing mass maps. These tests have demonstrated the self-consistency of our modeling framework. The correctness of our model has been further verified by comparing our predictions with observed number counts of shear-selected clusters as a function of both weak-lensing signal-to-noise ratio and richness of the optical counterparts. These validation tests have been carried out using a small subfield of the HSC-Y1 shear catalog so that we still respect the blinded analysis.
@@ Acknowledgements
We thank an anonymous referee for valuable comments and suggestions. K.-F.C. acknowledges support from the Taiwan Think Global Education Trust Scholarship and the Taiwan Ministry of Education’s Government Scholarship to Study Abroad. K.-F.C. thanks the computing resources provided by the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan which are maintained by Dr. Bau-Ching Hsieh and others. K.-F.C. thanks the hospitality of Adrian Liu at McGill University where this work is carried out in part. I-Non Chiu acknowledges the support from the National Science and Technology Council in Taiwan (Grant NSTC 111-2112-M-006-037-MY3) and the computing resources provided by the National Center for High-Performance Computing (NCHC) in Taiwan. YTL acknowledges supports by the grant NSTC 112-2112-M-001-061. This work was supported by JSPS KAKENHI Grant Numbers JP20H05856, JP22H01260, JP22K21349.
@@
This paper makes use of software developed for Vera C. Rubin Observatory. We thank the Rubin Observatory for making their code available as free software at http://pipelines.lsst.io/. Numerical calculations are performed through the Python packages numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020). Many of the cosmological and astrophysical calculations in this work rely on the routines wrapped up in colossus (Diemer, 2018) and astropy (Astropy Collaboration et al., 2022). Plots are made available thanks to matplotlib (Hunter, 2007) and seaborn (Waskom, 2021).
@@ Appendix A Confidence level of Cross-matched Optical Counterparts
Murata et al. (2019) obtained constraints on the parameter $(A,B,B_{z},C_{z},\sigma_{0},q,q_{z},p_{z})$ by fitting to the HSC-Y1 shear data assuming a pivot mass $M_{\mathrm{pivot}}=3\times 10^{14}\,h^{-1}M_{\odot}$ and redshift $z_{\mathrm{pivot}}=0.6$. Here, we adopt the constraint derived under the Planck Collaboration et al. (2020) cosmology $(A,B,B_{z},C_{z},\sigma_{0},q,q_{z},p_{z})=(3.36,0.83,-0.20,3.51,0.19,-0.02,0.23,1.26)$. We note that this constraint is obtained for the optical clusters in the CAMIRA catalog with $N_{\mathrm{mem}}\geq 15$ and $0.1\leq z_{\mathrm{cl}}\leq 1.0$. Therefore, we only cross-match to CAMIRA clusters within these ranges. For the other terms in Equation (34), $P\left(\nu|M,z\right)$ is obtained following the treatment in Sec. 4.3
@@
Figure 9: Same as Figure 8, number counts of weak lensing peaks’ optical counterparts as a function of their optical richness given the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived with different allowed range of scattering $\Delta_{s}$. These results are derived with $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ for HSC-Y3 after we unblind the cosmological analysis in Chiu et al. (2024). The green squares show the number counts of the closest CAMIRA counterpart for weak-lensing peaks in the entire HSC-Y3 footprint with a signal-to-noise ratio larger than $4.7$.